answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What are the odds of selecting a four of diamonds out of a deck of fiffty-two card?

The answer depends on the number of draws, whether or not hey are random, whether or not cards are returned to the deck before the next draw.

If only one card that is drawn at random, the probability is 1/52.

How do you predict probability?

"Probability" is not something that occurs in the future. It's the

numerical likelihood of something happening in the future.

You don't predict the probability. You calculate it.

How can you identify marble?

it is a heavy, solid rock with charcaterisitc white colouring

What is the standard notation to represent the probability of an event?

It is on a scale of 1 to 0 of a probability that an event will happen or it will not happen

What Key word do you associate with sequence of events and what rules of probability do you use?

The key word that I use is causality. However, you cannot use probability to determine causality. Even if two events are highly correlated, probability theory cannot tell whether event A is caused by event B, or event B is caused by event A, or that both are caused by some third event that is not even part of the study.

A class of 14 boys and 11 girls. What is the probability of selecting a girl 1st and then a girl 2nd and then a boy third.?

Assuming the choices are made randomly and that the chosen people are not returned to the class, the probability is 77/690 = 0.1116 approx.

What is the probability of a square?

The answer depends on what the experiment is. And since you have not bothered to share that crucial bit of information, I cannot provide a more useful answer.

What is greater 1sixth or 1 eighth?

The greater the denominator the smaller its value. Since eight is greater than six the value of 1eighth is lesser than 1 sixth.

Another way of saying,

1 eighth is nothing but dividing 1 into eight equal parts and 1 sixth is dividing into six equal parts. Obviously the 1 sixth part will be greater than one eighth. For eg. if you share a cake equally with eight friends, the slice be will smaller than when you share equally with six friends.

How would you find the probability that you would toss two heads in a row with a coin?

Here are two ways:

You could make a table:

. |H | T

---------

H| 1 0

T| 0 0

The row across the top is the first toss. The column is the second toss. The one with both Heads is indicated with a 1. There is 1 chance out of the 4 possible outcomes, so 1/4 = 0.25

Mathematically: Chance of first coin Heads = 0.5, chance of 2nd coin heads = 0.5; then multiply the two probabilities together, since they both have to happen: (0.5)*(0.5) = 0.25

What is the chance of living to the age of 1000?

Pretty unlikely as the oldest person who ever lived died at 122 years and 164 days.

After a breakup is a narcissist less likely to contact you if he or she believes you have figured them out?

If they know that you have figured them out, they will not want to be around you for fear that you will tell people the truth. What they really need is help for their problem.

What are the odds of rolling a pair with two dice?

1 in 6.

Basically dice 1 can roll any number

So the odds are that dice 2 has the same number as dice 1, or 1 in 6

What is the probability of rolling an even number on either die if two dice are rolled simultaneously?

There is a 1/2 chance of rolling an odd number on either die, so 1/2 X 1/2 = 1/4. 1/4 is the probability that both die will have odd numbers, so 3/4 is the probability that at least one is even.

Trending Questions
How many children is it possible to have at once? A couple has two female children What is the probability that their next child will be male? Do you agree or disagree with david hume's assesment that we can know the external world with neither certainty nor any probability? If you flip 10 pennies all at the same time, what is the probability that exactly 3 will come up heads and 7 will come up tails? What is the probability of sun rising in east? What is the probability that you do not pick a picture card? How many different ways can you arrange 20 photos in an array? Why don't all normal curves have the same standard deviation? When a standard windsock becomes fully inflated the windspeed equals or exceeds? If you have a spinner numbered 1 to 5 What is the probability of spinning a 2? What is the theoretical probability of spinning an even number on a spinner is 23. The spinner has 8 even-number sections. How many sections are on the spinner? What is the probability of rolling Two dice with a sum of 12? Which of the following best explains step 2 of the ORM six step process? What is the probability of getting a sum greater than 5 if you roll 2 dice? What does the letter f mean in probability? What is the probability that a three digit number chosen at random is dividable by 2? Which two groups would have had shaved heads at certain times throughout history? How many ways can you arrange the letters in the word FROG? The expected value is what kind of aspect of how probability distribution is characterized? The probability of rolling an even number when number cube is rolled twice?